# Geometric-Aware Low-Light Image and Video Enhancement via Depth Guidance

### v1

@@ Abstract
Low-Light Enhancement (LLE) is aimed at improving the quality of photos/videos captured under low-light conditions. It is worth noting that most existing LLE methods do not take advantage of geometric modeling. We believe that incorporating geometric information can enhance LLE performance, as it provides insights into the physical structure of the scene that influences illumination conditions. To address this, we propose a Geometry-Guided Low-Light Enhancement Refine Framework (GG-LLERF) designed to assist low-light enhancement models in learning improved features for LLE by integrating geometric priors into the feature representation space. In this paper, we employ depth priors as the geometric representation. Our approach focuses on the integration of depth priors into various LLE frameworks using a unified methodology. This methodology comprises two key novel modules. First, a depth-aware feature extraction module is designed to inject depth priors into the image representation. Then, Hierarchical Depth-Guided Feature Fusion Module (HDGFFM) is formulated with a cross-domain attention mechanism, which combines depth-aware features with the original image features within the LLE model. We conducted extensive experiments on public low-light image and video enhancement benchmarks. The results illustrate that our designed framework significantly enhances existing LLE methods.
Figure 1: We propose to enhance the performance of existing LLIE and LLVE methods via depth priors via our designed GG-LLERF. Our method consistently achieves the improvement target on different datasets/target networks with the same structure. “R.F.” denotes Retinexformer, “R.T.” represents Restormer.
1 Introduction
Low-light imaging is a common requirement in our daily lives, but it often results in poor quality images or videos due to inadequate illumination or limited exposure time [3]. In response to this issue, the Low-Light Enhancement (LLE) technique has been developed. Its objectives include reducing noise and artifacts, preserving edges and textures, and reproducing natural brightness and color [17, 28].
@@ 1 Introduction
One approach to improving enhancement performance involves the utilization of multi-modal maps as priors. For example, SKF [23] leverages semantic maps to optimize the feature space for low-light enhancement, while SMG [27] employs a generative framework to integrate edge information, thereby enhancing the initial appearance modeling for low-light scenarios. However, it’s worth noting that all of these employed priors operate at the 2D level. 2D-level information may not fully capture the geometric information of the entire 3D scene. The inclusion of 3D geometric information can be instrumental in determining the illumination distribution in the scene, thereby enhancing LLE tasks. This is especially important in video tasks. Having geometric information of each frame allows for a more accurate modeling of the scene’s geometric features.
@@
In this paper, we introduce a groundbreaking approach by learning the essential geometric priors within Low-Light Enhancement (LLE) frameworks. Subsequently, we develop an effective Geometry-Guided Low-Light Enhancement Refine Framework (GG-LLERF) tailored to incorporate these priors into the target frameworks $\mathcal{G}$, thereby surpassing their original performance limits. For our geometric representation, we harness depth priors, which offer valuable insights into the scene’s geometry. We choose depth priors because of the existence of highly effective open-world depth estimation models $\mathcal{D}$ that have been trained on large-scale datasets, e.g., DPT [13].
@@
Formulating a desired depth prior presents two primary challenges. First, achieving precise depth estimation directly from a low-light image is a demanding task. Existing depth estimation networks often struggle to provide ideal zero-shot predictions under various dark environments. Second, a significant domain gap exists between the depth scores and image content. Consequently, the direct fusion of depth and image features proves to be a complex and challenging endeavor.
@@
To address the first challenge, we introduce an encoder-decoder network denoted as $\mathcal{F}$, which predicts depth information from low-light images/frames $I_{l}$. We use the depth prediction from the corresponding normal-light images $I_{n}$ as the ground truth for supervision. This approach allows us to distill the depth prediction capability of $\mathcal{D}$ to achieve the desired depth estimation in low-light conditions. To overcome the second challenge, instead of directly fusing depth and image data, we propose a feature fusion method that involves the encoder of $\mathcal{G}$ and the depth-aware feature extraction module of $\mathcal{F}$ (with the extracted features denoted as $f_{g}$ and $f_{d}$, respectively). These features are extracted from $I_{l}$, making them more compatible and suitable for fusion. Furthermore, these depth-aware feature priors $f_{d}$ not only convey the geometrical information of the target scene but also encapsulate complementary image characteristics that can enhance the features $f_{g}$ extracted by $\mathcal{G}$.
@@
Once the depth-aware priors have been obtained and the fusion locations identified, the key challenge is how to effectively carry out the fusion process. To address this, we introduce the Hierarchical Depth-Guided Feature Fusion Module (HDGFFM) at various encoder stages of $\mathcal{G}$. The HDGFFM facilitates feature fusion between $f_{g}$ and $f_{d}$ through a cross-attention strategy, where the depth-aware features $f_{d}$ serve as the query vector, while the extracted feature $f_{g}$ functions as the key and value vectors in the transformer’s attention computation. Compared to self-attention using $f_{g}$ alone, this cross-attention strategy allows us to obtain depth-aware features that encompass information beyond the image content. These depth-aware features can then be harnessed to refine $f_{g}$ for enhancement, guided by the objectives of LLE. The proposed HDGFFM with its cross-attention strategy is adaptable to various LLE networks, making it a versatile and effective fusion method.
@@
Extensive experiments are conducted on public LLIE and LLVE datasets with various target networks. Experimental results demonstrate that our GG-LLERF can be combined with existing LLIE and LLVE frameworks to improve their performance on varying datasets. In summary, our contribution is three-fold.
  • We are the first to propose adopting the designed suitable depth-aware features as the prior for LLIE and LLVE tasks, and we formulate the corresponding Geometry-Guided Low-Light Enhancement Refine Framework.
  • HDGFFM is proposed to conduct feature fusion in an effective attention computation manner.
  • Extensive experiments are conducted on different datasets and networks, showing the effectiveness of our strategy.
2 Related Work
@@
Figure 2: The overall framework of GG-LLERF, which incorporates depth-aware features into the encoder of the target LLE framework (LLIE or LLVE) $\mathcal{G}$. The depth-aware features are extracted from the depth modeling network $\mathcal{F}$, which is supervised by a pre-trained depth estimator $\mathcal{D}$. The depth-aware features and image/frame features are fused via our proposed Hierarchical Depth-Guided Feature Fusion Module (HDGFFM) (Sec. 3.2), where Correlation-based Cross Attention is utilized (Sec. 3.3).
2.1 Low-light Image and Video Enhancement
@@
In recent years, there have been notable advancements in learning-based LLIE techniques [32, 24, 34, 8, 36, 37, 20, 10, 30, 7, 29], primarily emphasizing supervised approaches due to the availability of abundant image pairs for training. For example, MIRNet [31] adopts a multi-scale architecture to effectively capture and distill information at various levels. This results in improved image quality with enhanced brightness, contrast, and details, along with a reduction in noise. Fu et al. [5] implemented illumination augmentation on a pair of images, successfully achieving self-supervised Retinex decomposition learning. This innovative approach contributes to further advancements in LLIE techniques.
@@
Beyond LLIE, there is a rising need for LLVE solutions. Danai et al. [14] proposed a data synthesis mechanism that generates dynamic video pairs static datasets, formulating a LLVE framework. Wang et al. [19] introduced a multi-task model capable of simultaneously estimating noise and illumination, particularly effective for videos with severe noise conditions. Xu et al. [26] innovatively designed a parametric 3D filter tailored for enhancing and sharpening low-light videos, contributing to the latest developments in video enhancement techniques.
@@
While significant progress has been achieved in both LLIE and LLVE, there remains a challenge in obtaining accurate enhancement results in certain demanding scenarios. This difficulty arises from the highly ill-posed nature of directly recovering normal-light photos from these images. Consequently, the incorporation of suitable priors becomes imperative to address these challenges effectively.
2.2 Low-light Enhancement with Priors
@@
3 Method
In this section, we first provide an overview of our proposed strategy, which leverages geometrical priors (i.e., depth), in Sec. 3.1. Following the overview, we delve into the structural details of our proposed Hierarchical Depth-Guided Feature Fusion (HDGFFM) in Sec. 3.2. This includes discussing how to obtain the desired depth prior for fusion, where the fusion takes place, and the methods employed for conducting the fusion. The next section is dedicated to describing a vital component of our fusion procedure, known as Correlation-based Cross Attention for Fusion, and can be found in Sec. 3.3. Finally, we introduce the training pipeline of our strategy in Sec. 3.4.
3.1 Overview
@@
Depth estimation has been a long-standing task, and it has recently reached new heights with the development of large models. By training on extensive datasets that encompass diverse scenes and objects, current depth estimation models (e.g., DPT [13]) exhibit remarkable zero-shot performance on various images and videos. Consequently, we suggest distilling depth information from SOTA depth prediction networks and integrating it into the low-light enhancement task.
Implementation. In this paper, we focus on the supervised learning manner for LLE, where low-light data is denoted as $I_{l}$, and the normal-light data is represented as $I_{n}$. As shown in Fig., we set a lightweight depth estimation branch $\mathcal{F}$ for a given target LLE framework $\mathcal{G}$ to be improved.
$\mathcal{F}$ takes the input of $I_{d}$ and output $\hat{d}_{d}$, as
$\hat{d}_{d}=\mathcal{F}(I_{l};\theta_{f}),$ (1)
@@
We combine the extracted features from $\mathcal{G}$ and $\mathcal{F}$ using a cross-attention module $\mathcal{C}$. This process refines the features in $\mathcal{G}$ by incorporating depth information from $\mathcal{F}$, ultimately enhancing the final results, denoted as $\hat{I}_{n}$, as
$\hat{I}_{n}=\mathcal{G}(I_{d},\mathcal{F}_{e}(I_{l});\theta_{g}),$ (2)
where $\mathcal{F}_{e}(I_{l})$ denotes the extracted feature with $\mathcal{F}$. The objective can be represented as
$\hat{\theta}_{g},\hat{\theta}_{f}=argmin(\mathcal{L}(\hat{I}_{n},I_{n})+\mathcal{L}(\hat{d}_{d},d_{n})).$ (3)
@@
3.2 Hierarchical Depth-Guided Feature Fusion
What and where to fuse? Assuming we can obtain the desired depth prediction $\hat{d}_{d}$ from the low-light input data, it remains challenging to directly integrate $\hat{d}_{d}$ into the enhancement process of $\mathcal{G}$ due to the fundamental difference between depth and image domains. Therefore, we propose conducting the fusion in the deep feature space, specifically between the extracted features from $\mathcal{G}$ and $\mathcal{F}$, which we denote as $f_{g}$ and $f_{d}$, respectively. This approach effectively mitigates the challenges posed by the disparity between the two data sources during fusion, as both $f_{g}$ and $f_{d}$ are derived from the same input image $I_{d}$.
@@ 3.2 Hierarchical Depth-Guided Feature Fusion
$\hat{f}_{d,l}=\mathcal{E}_{l}(f_{d,l}).$ (4)
The depth-aware embedding modules serve the purpose of channel adjustment and do not involve information compression. Therefore, they are designed to be lightweight, consisting of just a single layer of linear convolution.
The fusion process is carried out through our proposed cross-attention mechanism, which will be elaborated on in the upcoming section.
Figure 3: The overview of “Correlation-based Cross Attention for Fusion”. The attention map is computed in both self-attention 5 and cross-attention between different domains 7.
3.3 Correlation-based Cross Attention for Fusion
Feature fusion strategy. Several strategies have been employed to fuse cross-domain features for low-light enhancement. In SMG [27], synthesized edge information serves as a condition for generating feature normalization and convolution parameters. However, this operation can be computationally expensive, especially when dealing with high-resolution images. On the other hand, SKF [23] computes the similarity between image and semantic features and utilizes this similarity to modulate the original image features, using query, key, and value features in the fusion as image, semantic, and image features. It’s worth noting that this strategy may not effectively incorporate new information, as the semantic features are primarily used to facilitate the computation of an attention matrix rather than being integrated with the image features themselves.
In this paper, our approach involves directly guiding the depth-aware features into the pathways of $\mathcal{G}$. Additionally, we integrate the results of the cross-attention computation with the outputs of the self-attention computation, based on a feature correlation measurement.
Fusion computation details. Suppose our goal is to fuse features in the $l$-th layer, denoted as $f_{g,l}$ and $\hat{f}_{d,l}$. Both of these feature maps have the same shape of $H\times W\times C$, where $H$ represents the feature height, $W$ the feature width, and $C$ the feature channel. To perform the feature computation, separate projection matrices for query, key, and value [15] are established for both $f_{g,l}$ and $\hat{f}_{d,l}$. For $f_{g,l}$, these projection matrices are labeled as $W_{q}$, $W_{k}$, and $W_{v}$, while for $\hat{f}_{d,l}$, only the key and query matrices are required, designated as $\hat{W}_{k}$ and $\hat{W}_{v}$, respectively.

### v2

@@ Abstract
Low-Light Enhancement (LLE) is aimed at improving the quality of photos/videos captured under low-light conditions. It is worth noting that most existing LLE methods do not take advantage of geometric modeling. We believe that incorporating geometric information can enhance LLE performance, as it provides insights into the physical structure of the scene that influences illumination conditions. To address this, we propose a Geometry-Guided Low-Light Enhancement Refine Framework (GG-LLERF) designed to assist low-light enhancement models in learning improved features by integrating geometric priors into the feature representation space. In this paper, we employ depth priors as the geometric representation. Our approach focuses on the integration of depth priors into various LLE frameworks using a unified methodology. This methodology comprises two key novel modules. First, a depth-aware feature extraction module is designed to inject depth priors into the image representation. Then, the Hierarchical Depth-Guided Feature Fusion Module (HDGFFM) is formulated with a cross-domain attention mechanism, which combines depth-aware features with the original image features within LLE models. We conducted extensive experiments on public low-light image and video enhancement benchmarks. The results illustrate that our framework significantly enhances existing LLE methods. The source code and pre-trained models are available at https://github.com/Estheryingqi/GG-LLERF.
Index Terms:
Low-Light Enhancement, Depth-aware Feature, Cross-domain Attention
Figure 1: We propose to enhance the performance of existing LLIE and LLVE methods via depth priors with our designed $G$-LLERF. Our method consistently achieves the improvement target on different datasets/networks with the same structure to extract depth-aware features and complete the fusion with original images/frames’ features.
I Introduction
Low-light imaging is a common requirement in our daily lives, but it often results in poor-quality images or videos due to inadequate illumination or limited exposure time [1]. In response to this issue, the Low-Light Enhancement (LLE) technique has been developed. Its objectives include reducing noise and artifacts, preserving edges and textures, and reproducing natural brightness and color [2, 3].
@@ I Introduction
One approach to improving enhancement performance involves the utilization of multi-modal maps as priors. For example, SKF [12] leverages semantic maps to optimize the feature space for low-light enhancement; SMG [13] employs a generative framework to integrate edge information, thereby enhancing the initial appearance modeling for low-light scenarios. However, it’s worth noting that all of these employed priors operate at the 2D level. 2D-level information may not fully capture the geometric information of the entire 3D scene. The inclusion of 3D geometric information can be instrumental in determining the illumination distribution in the scene, thereby enhancing LLE tasks. This is especially important in video tasks. Having geometric information of each frame allows for a more accurate modeling of the scene’s geometric features.
@@
In this paper, we introduce a groundbreaking approach by learning the essential geometric priors within Low-Light Enhancement frameworks. Subsequently, we develop an effective Geometry-Guided Low-Light Enhancement Refine Framework (GG-LLERF) tailored to incorporate these priors into the target frameworks $\mathcal{G}$, thereby surpassing their original performance limits. For our geometric representation, we harness depth priors, which offer valuable insights into the scene’s geometry. We choose depth priors because of the existence of highly effective open-world depth estimation models $\mathcal{D}$ that have been trained on large-scale datasets, e.g., DPT [14].
@@
Formulating a desired depth prior presents two primary challenges. First, achieving precise depth estimation directly from a low-light image is a demanding task. Existing released depth estimation networks $\mathcal{D}$ often struggle to provide ideal zero-shot predictions under various dark environments. Second, a significant domain gap exists between the depth scores and image content. Consequently, the direct fusion of depth and image features proves to be a challenging endeavor.
@@
To address the first challenge, we introduce an encoder-decoder network denoted as $\mathcal{F}$, which predicts depth information from low-light images/frames $I_{d}$. We use the depth prediction from the corresponding normal-light images $I_{n}$ via $\mathcal{D}$ as the ground truth for supervision. This approach allows us to distill the depth prediction capability of $\mathcal{D}$ to achieve the desired depth estimation in low-light conditions (the features of pre-trained network $\mathcal{D}$ will not be employed during training and inference). To overcome the second challenge, instead of directly fusing depth and image data, we propose a feature fusion method that involves the encoder of $\mathcal{G}$ and the depth-aware feature encoder of $\mathcal{F}$ (with the extracted features denoted as $f_{g}$ and $f_{d}$, respectively). These features are extracted from $I_{d}$, making them more compatible and suitable for fusion. Furthermore, these depth-aware feature priors $f_{d}$ not only convey the geometrical information of the target scene but also encapsulate complementary image characteristics that can enhance the features $f_{g}$ extracted by $\mathcal{G}$.
@@
Once the depth-aware priors have been obtained and the fusion locations identified, the critical challenge is how to effectively carry out the fusion process. To address this, we introduce the Hierarchical Depth-Guided Feature Fusion Module (HDGFFM) at various encoder stages of $\mathcal{G}$. The depth-aware features represent the information of texture, edge, depth, and distance. HDGFFM facilitates feature fusion between $f_{g}$ and $f_{d}$ through a cross-attention strategy, where the depth-aware features $f_{d}$ serve as the query vector, while the extracted feature $f_{g}$ functions as the key and value vectors in the transformer’s attention computation. Compared to self-attention using $f_{g}$ alone, this cross-attention strategy allows us to obtain depth-aware features that encompass information beyond the image content and clearly distinguish the boundaries of different objects. These depth-aware features can then be harnessed to refine $f_{g}$ for enhancement, guided by the objectives of LLE. The proposed HDGFFM, with its cross-attention strategy, is adaptable to various LLE networks, making it a versatile and effective fusion method.
@@
Extensive experiments are conducted on public LLIE and LLVE datasets with various target networks. Experimental results demonstrate that our GG-LLERF can be combined with existing LLIE and LLVE frameworks to improve their performance on varying datasets. The improvement is more significant compared to other methods using different priors, such as SKF [12]. In summary, our contribution is three-fold.
  • To the best of our knowledge, we are the first to propose using specially designed depth-aware features from input low-light images as priors for LLIE and LLVE tasks. While some previous methods rely on depth features extracted from the depth map, we uniquely assert the rationality of depth-aware features. The formulation strategy of depth-aware features is one of our key contributions.
  • We have designed the overall structure of the Geometry-Guided Low-Light Enhancement Refine Framework, including the points at which feature fusion is conducted and the type of depth priors to be fused. This strategy is plug-and-play, allowing it to be seamlessly integrated into any low-light enhancement network.
  • We propose the HDGFFM to perform feature fusion between depth-aware and image content features using a novel attention computation method. Compared to previous strategies, our approach integrates both self-attention and cross-attention between image and depth-aware features, along with the correlation modeling.
II Related Work
@@
Figure 2: The overall framework of $G$-LLERF, which incorporates depth-aware features into the encoder of the target LLE framework (LLIE or LLVE) $\mathcal{G}$. The depth-aware features are extracted from the depth modeling network $\mathcal{F}$, which is supervised by a pre-trained depth estimator $\mathcal{D}$. $\mathcal{F}$ is trained from scratch, not pre-trained. The pre-trained depth estimator $\mathcal{D}$, e.g., DPT, is just applied to the normal-light data and provides the pseudo ground truth of depth for training. $\mathcal{D}$ is frozen, and its features are not utilized. Moreover, the depth-aware features (from the encoder of $\mathcal{F}$) and image/video features are fused via our proposed Hierarchical Depth-Guided Feature Fusion Module (HDGFFM) (Sec. III-B), where Correlation-based Cross Attention is utilized (Sec. III-C). In contrast, the hierarchical features from the decoder of $\mathcal{F}$ are not employed during inference. Note that $\mathcal{G}$ can take either an image or a video as input, as it can employ both LLIE and LLVE structures. $\mathcal{F}$ is also capable of estimating depth for both images and videos, as its encoder shares the same input as that of $\mathcal{G}$, enabling effective feature fusion in HDGFFM.
II-A Low-light Image and Video Enhancement
@@
In recent years, there have been notable advancements in learning-based LLIE techniques [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25], primarily emphasizing supervised approaches due to the availability of abundant image pairs for training. For example, MIRNet [26] adopts a multi-scale architecture to effectively capture and distill information at various levels. This results in improved image quality with enhanced brightness, contrast, and details, along with a reduction in noise. Fu et al. [6] implemented illumination augmentation on a pair of images, successfully achieving self-supervised Retinex decomposition learning. This innovative approach contributes to further advancements in LLIE techniques. Zhang et al. [27] focused on using semantic segmentation priors for low-light image enhancement, enabling locally controllable enhancement. Depth information was used auxiliary to the semantic priors. Moreover, they directly obtained depth features from predicted depth maps during inference. Wu et al. [12] incorporated semantic information, primarily utilizing a pre-trained frozen segmentation network to obtain feature priors from low-light images, whereas the pre-trained frozen model may suffer from domain shift. Recently, diffusion-based models also appear [28, 29].
@@
Beyond LLIE, there is a rising need for LLVE solutions. Danai et al. [30] proposed a data synthesis mechanism that generates dynamic video pairs static datasets, formulating a LLVE framework. Wang et al. [7] introduced a multi-task model capable of simultaneously estimating noise and illumination, particularly effective for videos with severe noise conditions. Xu et al. [8] innovatively designed a parametric 3D filter tailored for enhancing and sharpening low-light videos, while cannot be utilized as a play-and-plug strategy.
@@
While significant progress has been achieved in both LLIE and LLVE, there remains a challenge in obtaining accurate enhancement results in certain demanding scenarios. This difficulty arises from the highly ill-posed nature of directly recovering normal-light photos from these images. The ill-posed issue refers to the fact that the ground truth for an input image can vary, but only a single ground truth is provided in the training dataset. Moreover, degradation processes in real-world scenarios are often complex and unknown, leading to irreversible damage to images. As a result, various plausible outcomes can arise from the low-quality input during low-light image enhancement. Meanwhile, a representative phenomenon for the ill-posed issue occurs where some enhancement results closely approximate the target, while others do not. Adding priors helps to impose additional constraints on the enhancement process, reducing the range of plausible outcomes. Consequently, the output is closer to the distribution of the ground truth. Therefore, the incorporation of suitable priors becomes imperative to address these challenges effectively.
II-B Low-light Enhancement with Priors
@@
II-C Differences with Current Depth-aware Restoration
Several frameworks have been developed to incorporate depth awareness into image restoration tasks. Bao et al. [31] proposed a video frame interpolation method that explicitly detects occlusions by exploiting depth information. Cheng et al. [32] utilized depth information from a pre-trained monocular depth estimation model to restore high-resolution (HR) images in a self-supervised manner. Liet al. [33] introduced a deep neural convolutional network that leverages depth maps for dynamic scene deblurring. Lee et al. [34] proposed a geometry-aware deep video deblurring method. However, none of these methods were designed specifically for the LLE task, and their strategies for fusing depth and image information differ from our approach. They either utilize depth information as supervision or directly fuse features extracted from depth maps and images. In contrast, our paper presents a novel framework that leverages depth information by distilling depth prediction ability for low-light images (with the depth estimator trained from scratch), and fusing image content features with depth-aware image features that are not extracted from the depth map.
III Method
In this section, we first provide an overview of our proposed strategy, which leverages geometrical priors (i.e., depth), in Sec. III-A. Following the overview, we delve into the structural details of our proposed Hierarchical Depth-Guided Feature Fusion (HDGFFM) in Sec. III-B. This includes discussing how to obtain the desired depth prior for fusion, where the fusion takes place, and the methods employed for conducting the fusion. The following section is dedicated to describe a vital component of our fusion procedure, known as Correlation-based Cross Attention for Fusion, and can be found in Sec. III-C. Finally, we introduce the training pipeline of our strategy in Sec. III-D.
III-A Overview
@@
Depth estimation has been a long-standing task, and it has recently reached new heights with the development of large models. By training on extensive datasets that encompass diverse scenes and objects, current depth estimation models (e.g., DPT [14]) exhibit remarkable zero-shot performance on various images and videos. Consequently, we suggest distilling depth information from SOTA monocular depth prediction networks and integrating it into the LLE task.
Implementation. In this paper, we focus on the supervised learning manner for LLE, where low-light data is denoted as $I_{d}$, and the normal-light data is represented as $I_{n}$. As shown in Fig. 2, we set a lightweight depth estimation branch $\mathcal{F}$ for a given target LLE framework $\mathcal{G}$ to be improved. $\mathcal{F}$ takes the input of $I_{d}$ and output $\hat{d}_{d}$, as
$\small\hat{d}_{d}=\mathcal{F}(I_{d};\theta_{f}),$ (1)
@@
We combine the extracted features from the encoder of $\mathcal{G}$ and $\mathcal{F}$ using a cross-attention module $\mathcal{C}$. This process refines the features in the encoder of $\mathcal{G}$ by incorporating depth information from $\mathcal{F}$, ultimately enhancing the final results, denoted as $\hat{I}_{n}$, as
$\small\hat{I}_{n}=\mathcal{G}(I_{d},\mathcal{F}_{e}(I_{d});\theta_{g}),$ (2)
where $\mathcal{F}_{e}(I_{d})$ denotes the extracted feature with $\mathcal{F}$. The objective can be represented as
$\small\{\hat{\theta}_{g},\hat{\theta}_{f}\}=argmin(\mathcal{L}(\hat{I}_{n},I_{n})+\mathcal{L}(\hat{d}_{d},d_{n})).$ (3)
@@
III-B Hierarchical Depth-Guided Feature Fusion
How to fuse? Assuming we can obtain the desired depth prediction $\hat{d}_{d}$ from the low-light input data, it remains challenging to directly integrate $\hat{d}_{d}$ into the enhancement process of $\mathcal{G}$ due to the fundamental difference between depth and image domains. Therefore, we propose conducting the fusion in the deep feature space, specifically between the extracted features from $\mathcal{G}$ and $\mathcal{F}$, which we denote as $f_{g}$ and $f_{d}$, respectively. This approach effectively mitigates the challenges posed by the disparity between the two data sources during fusion, as both $f_{g}$ and $f_{d}$ are derived from the same input image $I_{d}$.
@@ III-B Hierarchical Depth-Guided Feature Fusion
$\small\hat{f}_{d,l}=\mathcal{E}_{l}(f_{d,l}).$ (4)
The depth-aware embedding modules adjust channels without information compression. Therefore, they are designed to be lightweight, consisting of a single layer of linear convolution.
The fusion process is carried out through our cross-attention mechanism, which will be elaborated on in the next section.
Figure 3: The overview of “Correlation-based Cross Attention for Fusion”. The attention map is computed in both self-attention (Eq 5) and cross-attention between different domains (Eq 7).
III-C Correlation-based Cross Attention for Fusion
Feature fusion strategy. Several strategies have been employed to fuse cross-domain features for low-light enhancement. In SMG [13], synthesized edge information serves as a condition for generating feature normalization and convolution parameters. However, this operation can be computationally expensive, especially when dealing with high-resolution images. On the other hand, SKF [12] computes the similarity between image and semantic features and utilizes this similarity to modulate the original image features, i.e., they use image, semantic, and image features as query, key, and value features in the fusion. It’s worth noting that this strategy may not effectively incorporate new information, as the semantic features are primarily used to facilitate the computation of an attention matrix rather than being integrated with the image features themselves. And our method’s superiority to these methods are shown in Table. VI.
In this paper, as shown in Fig. 3, our approach involves directly guiding the depth-aware features into the pathways of $\mathcal{G}$. Additionally, we integrate the results of the cross-attention computation with the outputs of the self-attention computation, based on a feature correlation measurement.
Fusion computation details. Suppose our goal is to fuse features in the $l$-th layer, denoted as $f_{g,l}$ and $\hat{f}_{d,l}$. Both of these feature maps have the same shape of $H\times W\times C$, where $H$ represents the feature height, $W$ the feature width, and $C$ the feature channel. To perform the feature computation, projection matrices for query, key, and value [37] are established for both $f_{g,l}$ and $\hat{f}_{d,l}$. For $f_{g,l}$, these projection matrices are labeled as $W_{q}$, $W_{k}$, and $W_{v}$, while for $\hat{f}_{d,l}$, only the key and query matrices are required, designated as $\hat{W}_{k}$ and $\hat{W}_{v}$, respectively.